OpenStudy (anonymous):

21 is 35% of what number? A.21=0.35 x x;62.4 B.x=21x 0.3k;7.4 C.0.35=21x x;40.8 D.21=0.35 x x;60

OpenStudy (anonymous):

60

OpenStudy (anonymous):

35% x = 21 -> x = 21/ .35 = 60

OpenStudy (anonymous):

ZachKetchup, you must be working with proportions in your class! Again, set it up so that 21 out of some number N is 35 percent. 21/N=35/100. Cross multiply to get 21*100=35*N. Solve for N to get (21*100)/35=N. So N=60 ; D. Since it looks like the answers want you to represent the percentage as a decimal, you can change all of the 35/100 or that fraction to be .35 instead. That should give you the same result though. If we say 21/N=.35, then we can multiply by N to both sides, 21=.35*N. Then divide. 21/.35=N. That will still yield 60, and D.